# REA-RL: Reflection-Aware Online Reinforcement Learning for Efficient Large Reasoning Models

### v1

@@ Abstract
Large Reasoning Models (LRMs) demonstrate strong performance in complex tasks but often face the challenge of overthinking, leading to substantially high inference costs. Existing approaches synthesize shorter reasoning responses for LRMs to learn, but are inefficient for online usage due to the time-consuming data generation and filtering processes. Meanwhile, online reinforcement learning mainly adopts a length reward to encourage short reasoning responses, but tends to lose the reflection ability and harm the performance. To address these issues, we propose REA-RL, which introduces a small reflection model for efficient scaling in online training, offering both parallel sampling and sequential revision. Besides, a reflection reward is designed to further prevent LRMs from favoring short yet non-reflective responses. Experiments show that both methods maintain or enhance performance while significantly improving inference efficiency. Their combination achieves a good balance between performance and efficiency, reducing inference costs by 35% without compromising performance. Further analysis demonstrates that our methods are effective by maintaining reflection frequency for hard problems while appropriately reducing it for simpler ones without losing reflection ability. Codes are available at https://github.com/hexuandeng/REA-RL.
Figure 1: Overthinking and non-reflective cases from GSM8k. The left column shows the output of DeepSeek-R1-Distill-Qwen-7B, which reflects eight times before finishing generation. The middle column presents the output after online RL training using length rewards, which only spends 103 tokens in “think” part and no reflection, where an error occurs (underlined). The right column shows our output, which uses a similar budget to R1-7B in reasoning but only performs a single reflection.
@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated impressive performance in downstream applications [16, 38, 18]. Their human-like deliberation and insightful self-reflection ability facilitate thorough question consideration and verification, thereby improving performance on complex reasoning tasks [10, 8]. However, this often leads to excessive reasoning with minimal performance benefit, i.e., overthinking, substantially increasing the inference cost [5, 31].
@@
Prior research [28, 39, 11] attempts to generate shorter reasoning responses for supervised fine-tuning (SFT) or reinforcement learning (RL) to encourage concise response generation. However, this method relies on static datasets, the distribution of which may deviate from the trained model, especially later in training, leading to suboptimal results [37]. Furthermore, the time-consuming data generation and filtering processes severely limit its feasibility as an online data generation solution.
@@
Towards these problems, another line of work [24, 35, 2] employs online reinforcement learning. To enhance the efficiency of online data generation, multiple reasoning paths are sampled in parallel for a single query, with accuracy and length rewards used to encourage correct and concise responses [46, 43]. However, this self-iterative training paradigm can lead to unpredictable behavior. As illustrated in Figure 1, it can cause the LRM to completely lose its reflection capability, reverting to a naive chain-of-thought style, resulting in suboptimal performance on complex reasoning tasks.
@@
In this paper, we propose REA-RL, a reflection-aware reinforcement learning approach that integrates the strengths of aforementioned ones. First, we introduce a small reflection model into the online RL process to identify the first reflection position in a sampled response that is very likely to derive the final answer, and truncate the response to a shorter one for optimization. It enables us to perform both parallel sampling and sequential revisions, which has been demonstrated to achieve computationally optimal test-time scaling during inference [36]. Second, we introduce a reflection reward to prevent the models from favoring short yet non-reflective responses. The reward is calculated based on the density of keywords (e.g., “wait”, “alternatively”) that signal reflective thinking in the response.
@@
Experimental results reveal that employing only the length reward in online RL leads to a substantial performance degradation. In contrast, both the reflection model and the reflection reward significantly improve performance and prevent non-reflective behavior. The reflection reward is better at maintaining performance, while the reflection model achieves higher efficiency. Combining both approaches achieves a 35% response shortening on all datasets without performance reduction, and 44% on easier datasets. Further analysis demonstrates that these improvements stem from maintaining the reflection tendency on difficult questions and appropriately reducing reflection on simple questions without losing reflection ability. Our contributions are summarized as follows:
@@
  • We design a reflection reward to prevent non-reflective behavior in online RL and significantly enhance model performance compared to using length reward only.
  • Results demonstrate that each method reduces inference cost while preserving model performance. Combining both methods achieves a 35% response shortening without performance degradation, reducing the reflection frequency for simple questions while maintaining it for difficult ones.
@@ 2 Related Work
Large reasoning models.
System 1 large language models (LLMs) have shown remarkable capabilities across diverse tasks [1, 21, 42]. However, complex reasoning tasks like advanced mathematics and formal logic necessitate more structured analysis [40, 9]. To address this, LRMs have emerged [16, 10, 8]. These models explicitly generate intermediate reasoning steps, i.e., test-time scaling [4, 27], before providing a final answer, incorporating self-reflection and error correction to enhance their performance on complex reasoning tasks [26, 20]. However, this enhanced capability comes at the cost of slower and more verbose reasoning processes [41, 31].
@@
Training on revised responses with less overthinking via SFT and RL enhances reasoning efficiency. Several methods employ parallel sampling. Munkhbat et al. [28] construct concise reasoning via best-of-N sampling and few-shot conditioning. Yu et al. [44] skip reasoning for high-confidence samples when facing perturbation. Other methods utilize stronger models to generate revisions. C3ot [17] employs GPT-4 [1] to compress reasoning while preserving key information. Chen et al. [5] use strong LLMs to detect overthinking and generate preference data. Han et al. [11] search for the optimal token budget and generate corresponding length responses using multiple strong LLMs. Still others rely on heuristic algorithms and time-consuming post-verification and filtering. SPIRIT-FT [7] identifies crucial steps using perplexity difference after removing the step. LM-skip [23] stimulates step-skipping behavior by iteratively refining models to generate shorter and accurate reasoning paths. TokenSkip [39] achieves fine-grained compression by omitting less important tokens in CoT outputs detected by LLMLingua-2 [29]. However, these methods often suffer from low efficiency due to complex generation and filtering processes, leading to more than double the generation cost compared to naive sampling, rendering their efficiency too low for online settings.
@@
DeepSeek R1 [10] achieves promising results using only accuracy and format rewards with online RL, demonstrating its effectiveness. Several methods also use RL to enhance the efficiency of LLMs with length rewards, often normalized by a baseline budget. O1-Pruner [24] estimates the budget from a reference model, while Arora and Zanette [3] do so from per-prompt groups. Kimi K1.5 [8] normalizes length with the minimum and maximum lengths of generations, while GRPO-LEAD [46] bases it on the length distribution. Other approaches predict the required maximum length budget. DAST [35] estimates the budget based on problem difficulty. Yeo et al. [43] propose a cosine reward that incentivizes meaningful reasoning while only penalizing excessive length. L1 [2] expects the model to follow the maximum length limit in the prompt and designs a target-aware reward. Finally, Liu et al. [22] collect preference pairs and construct rewards based on relative length. However, most methods depend solely on multiple samplings of a query, which may cause unpredictable behavior in RL. To address this, we provide online revision and refined reward design, better guiding the model’s optimization direction.
@@ 3 Overthinking Detection
Previous works [8, 5, 31] observe that LRMs like QwQ-32B-Preview [38] and DeepSeek-R1 [10] tend to generate more solution rounds for simple math problems, allocating excessive computational resources with limited utility. We extend this analysis to smaller LRMs, and propose an effective detection method that does not necessitate the use of powerful closed-source LLMs for detection.
@@ 3.1 Auto-Detection of Overthinking
As illustrated in Figure 1, we observe a similar phenomenon in Deepseek-R1-Distill-Qwen-7B (R1-7B): the model tends to engage in excessive reflection after completing reasoning and obtaining the correct answer, leading to overthinking. In preliminary experiments, we find that weaker LLMs struggle to extract the boundaries of each reflection. However, we note that the conclusion of both reasoning and reflection is often a restatement of the answer. Therefore, we prompt LLMs to determine if each part of the response contains the correct answer, which is easier for LLMs. The thought process preceding the first occurrence of the correct answer is considered effective reasoning, while each subsequent appearance is an additional reflection, i.e., overthinking.
@@
We employ Qwen2.5-32B-Instruct [33] to identify these positions within the “think” part of LRMs. Specifically, we segment the think part into smaller chunks and provide the input question along with all these chunks as input, prompting the model to determine whether each chunk contains the correct answer. Subsequently, to ensure accuracy, we apply a further filtering step: chunks containing the answer are verified again one by one with the LLM, and only those confirmed as positive in both detections are considered correct. We design prompts with and without the inclusion of the gold answer as input, detailed in Appendix A.
@@
The tokens after the first correct answer are identified as overthinking and subsequently removed. Following this, we generate the revision by forcibly terminating the “think” part and compelling the LRM to continue generating the final answer, prefixed with "**Final Answer:**", which yields a revision without overthinking. For verification, the revision is deemed correct if the LRM can accurately generate the final answer. Finally, to prevent excessively long generations from skewing the average length, we limit the generation budget to 16k tokens.
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average
@@
Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 60.15 1007549 48.33 10010460 77.01 1005478
Fixed Trunc (90) 89.39 80.51 91.40 83.51 81.30 83.20 85.00 85.68 59.41 86.77 44.17 88.72 75.11 84.73
Fixed Trunc (80) 88.25 72.47 89.60 75.52 79.22 75.22 78.44 77.67 58.67 77.92 37.08 80.09 71.88 76.48
Fixed Trunc (70) 87.26 64.73 86.00 66.76 75.58 66.61 73.12 68.99 53.19 69.10 33.33 71.02 68.08 67.87
Model Revise 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 60.74 83.06 50.83 93.59 77.30 76.88
 + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 60.15 73.48 51.67 92.31 76.54 67.52
Table 1: Performance and efficiency after model revision. “Original” denotes the R1-7B performance before revision. “Acc” represents the response accuracy, and “TR” represents the efficiency compared to the original ones, calculated by the token ratio compared to the “Original”. Therefore, the TR in the first row is 100, representing no revision, and its subscript indicates the corresponding average token length. “Average” is the macro-average across datasets. The best and second-best results are bold and underline, respectively. Abbreviations of methods are defined in §3.2.
@@ 3.2 Efficiency of Auto-Detection
“Model Revise” and “+Gold” represent revision using our method without and with the gold answer as input, respectively. Our baseline for comparison is “Fixed Trunc ($G$)”, which denotes fixed truncation of the original generated response, retaining $n\%$ of the thinking tokens and truncating the response at the nearest newline. Subsequently, similar to response revision, the “think” part is terminated, and the LRM is compelled to generate the final answer.
Evaluation dataset.
We follow Yang et al. [42] to evaluate our approach on six math test sets, ordered by difficulty: GSM8K [6], 8.5K grade school math problems; MATH500 [13], 500 challenging high school competition problems; Gaokao23 [19], English-translated math questions from the 2023 Chinese Gaokao exam; Amc23111https://huggingface.co/datasets/AI-MO/aimo-validation-amc, 2023 American Mathematics Competitions; Olympiad [12], Olympiad-level math and science questions; and Aime24222https://huggingface.co/datasets/AI-MO/aimo-validation-aime, 2024 American Invitational Mathematics Examination. We employ the math validator provided by rStar [9]. Considering the limited size of Amc23 and Aime24, we sample 8 paths for each question to mitigate randomness.
@@
Results before and after revision are in Table 1. Without the gold answer as input, we can automatically remove 23% of tokens without performance degradation. When provided, 32% of tokens are removed with a minor performance decrease. However, fixed truncation shows a considerable performance decline with increasing truncation ratios. These results demonstrate the effectiveness of our detection method and underscore the severity of the overthinking problem.
Greater overthinking on simpler problems.
Across the two easiest, two medium, and two hardest math datasets, the overthinking tokens our method detects decrease progressively, at 38%, 31%, and 14%, respectively. This indicates that the overthinking problem is more prevalent in simpler datasets and less prevalent in more challenging ones.
@@ 4 Reflection-Aware Online Reinforcement Learning
Figure 2: Workflow of REA-RL. The upper section illustrates the baseline GRPO process, which parallel sampling $G$ completions $\{s_{1},...,s_{G}\}$, and optimize using an accuracy reward and a length reward ($\text{R}_{\text{Len}}$). The bottom section shows our reflection model ($\text{M}_{\text{Reflect}}$) performing sequential revision. From the $G$ generated completions, the reflection model identifies and truncates overthinking tokens (red segment), keeps only the preceding segments (yellow segment), and then lets the policy model ($\text{M}_{\text{Policy}}$) complete the answer generation (blue segment), resulting in the revisions $\{s_{1}^{r},...,s_{G}^{r}\}$, with cases shown in Appendix B.1. We also introduce a reflection reward ($\text{R}_{\text{Reflect}}$) and a refined length reward ($\text{R}_{\text{RLen}}$). Finally, the advantages $\{a_{1},...,a_{G},a_{1}^{r},...,a_{G}^{r}\}$ are calculated based on both the initial and revised completions, and used to train the policy model.
To maintain model performance while reducing inference cost, we employ online RL to enhance the distributional alignment between the data and the model (§4.1). Furthermore, to enable both parallel sampling and sequential revision for online RL data generation, thereby facilitating better scaling and guiding the direction of online RL optimization, we introduce an efficient reflection model that provides revisions online (§4.2). Finally, we refine the reward to include a reflection reward to penalize non-reflective behavior (§4.3). The workflow is illustrated in Figure 2.

### v2

# arXiv:2505.19862v2[cs.CL] 27 Feb 2026

@@ Abstract
Large Reasoning Models (LRMs) demonstrate strong performance in complex tasks but often face the challenge of overthinking, leading to substantially high inference costs. Existing approaches synthesize shorter reasoning responses for LRMs to learn, but are inefficient for online usage due to the time-consuming data generation and filtering processes. Meanwhile, online reinforcement learning mainly adopts a length reward to encourage short reasoning responses, but it tends to lose reflection ability and harm performance. To address these issues, we propose REA-RL, which introduces a small reflection model for efficient scaling in online training, offering both parallel sampling and sequential revision. Besides, a reflection reward is designed to further prevent LRMs from favoring short yet non-reflective responses. Experiments show that both methods maintain or enhance performance while significantly improving inference efficiency. Their combination achieves a good balance between performance and efficiency, reducing inference costs by 36% without compromising performance. Further analysis demonstrates that our methods are effective by maintaining reflection frequency for hard problems while appropriately reducing it for easier ones without losing reflection ability. Code is available at https://github.com/hexuandeng/REA-RL.
@@ 1 Introduction
Large Reasoning Models (LRMs) have demonstrated impressive performance in downstream applications (Jaech et al., 2024; Qwen, 2025; Kilpatrick, 2025). Their human-like deliberation and insightful self-reflection ability facilitate thorough question consideration and verification, thereby improving performance on complex reasoning tasks (Guo et al., 2025; Du et al., 2025). However, this often leads to excessive reasoning with minimal performance benefits, i.e., overthinking, substantially increasing the inference cost (Xu et al., 2025; Chen et al., 2025; Qu et al., 2025a).
Figure 1: Overthinking and non-reflective cases from GSM8k. The left shows the output of DeepSeek-R1-Distill-Qwen-7B (R1-7B), which reflects eight times before finishing generation. The middle presents the output after online RL training using length rewards, which only spends 103 tokens in “think” part and no reflection, where an error occurs (underlined). The right shows the output of our method, which uses a similar budget to R1-7B in reasoning but only performs a single reflection.
@@
Prior research (Munkhbat et al., 2025; Xia et al., 2025; Han et al., 2025) attempts to generate shorter reasoning responses for supervised fine-tuning (SFT) or reinforcement learning (RL) to encourage concise response generation. However, this method relies on static datasets, the distribution of which may deviate from the trained model, especially later in training, leading to suboptimal results (Tang et al., 2024). Furthermore, the time-consuming data generation and filtering processes severely limit its feasibility as an online data generation solution.
@@
To address these problems, another line of work (Luo et al., 2025a; Shen et al., 2025; Aggarwal and Welleck, 2025; Liang et al., 2025) employs online reinforcement learning. To enhance the efficiency of online data generation, multiple reasoning paths are sampled in parallel for a single query, with accuracy and length rewards used to encourage correct and concise responses (Zhang and Zuo, 2025; Yeo et al., 2025). However, this self-iterative training paradigm can lead to unpredictable behavior. As illustrated in Figure 1, it can cause the LRM to completely lose its reflection capability, reverting to a naive chain-of-thought style, resulting in suboptimal performance on complex reasoning tasks.
@@
In this paper, we propose REA-RL, a reflection-aware RL framework that integrates the strengths of the aforementioned ones. First, we introduce a small reflection model into the online RL process to identify the first reflection position in a sampled response that is very likely to lead to the final answer, and truncate the response to a shorter revision for optimization. It enables us to perform both parallel sampling and sequential revision, which has been demonstrated to achieve computationally optimal test-time scaling during inference (Snell et al., 2024). Second, we introduce a reflection reward to prevent the models from favoring short yet non-reflective responses. The reward is calculated based on the density of keywords (e.g., “wait”, “but”) that signal reflective thinking in the response.
@@
Results reveal that employing only the length reward leads to substantial performance degradation. In contrast, both the reflection model and the reflection reward improve performance and prevent non-reflective behavior. The reflection reward is better at maintaining performance, while the reflection model achieves higher efficiency. Combining the approaches achieves a 36% response shortening across all datasets without performance reduction. Further analysis demonstrates these improvements stem from maintaining the reflection tendency on difficult questions and appropriately reducing reflection on easy questions without losing reflection ability. Our contributions are:
@@
  • We design a reflection reward to prevent non-reflective behavior in online RL and significantly enhance model performance compared to using only the length reward.
  • Results demonstrate that each method reduces inference cost while preserving model performance. Combining both methods achieves a 36% response shortening without performance degradation.
@@ 2 Related Work
Training on revised responses with less overthinking via SFT and RL enhances reasoning efficiency. Several methods employ parallel sampling. Munkhbat et al. (2025) construct concise reasoning via best-of-N sampling. Yu et al. (2024) skip reasoning for high-confidence samples, while NoThink (Ma et al., 2025) forces LRMs to skip reasoning for all samples. Other methods utilize stronger models to generate shorter revisions (Kang et al., 2025; Chen et al., 2025; Han et al., 2025; Wen et al., 2025). Still others rely on heuristic algorithms and time-consuming post-verification. SPIRIT-FT (Cui et al., 2025) identifies crucial steps using perplexity. LM-skip (Liu et al., 2024c) stimulates step-skipping behavior by iteratively refining. TokenSkip (Xia et al., 2025) omits less important tokens detected by LLMLingua-2 (Pan et al., 2024). However, their efficiency is often too low for online settings due to complex generation and filtering processes, which incur more than double the generation cost compared to naive sampling.
@@
DeepSeek-R1 (Guo et al., 2025) achieves promising results using RL with verifiable rewards (RLVR). Several methods further introduce length rewards to enhance efficiency, often normalized by a baseline budget. O1-Pruner (Luo et al., 2025a) estimates the budget from a reference model, Liu et al. (2024a) from pairs, while Arora and Zanette (2025) from groups. ShorterBetter (Yi et al., 2025) estimates using shortest correct response, while Arora and Zanette (2025) uses Leave One Out estimator. Kimi K1.5 (Du et al., 2025) normalizes length with the minimum and maximum lengths of generations, while GRPO-LEAD (Zhang and Zuo, 2025) bases it on the length distribution. Other approaches predict the required maximum length budget. DAST (Shen et al., 2025) estimates the budget based on problem difficulty. Yeo et al. (2025) propose a cosine reward that only penalizes excessive length. L1 (Aggarwal and Welleck, 2025) rewards the model for following the length limit in prompts. However, most methods depend solely on multiple samplings of a query, which may cause unpredictable behavior in RL. To address this, we provide online revision and refined reward design, better guiding the model’s optimization direction.
@@ 3 Overthinking Detection
Previous works (Du et al., 2025; Chen et al., 2025; Xu et al., 2025; Qu et al., 2025a) observe that LRMs like QwQ-32B-Preview (Qwen, 2025) and DeepSeek-R1 (Guo et al., 2025) tend to generate more solution rounds for easy math problems, allocating excessive computational resources with limited utility. We extend this analysis to smaller LRMs, and propose an effective detection method that does not necessitate the use of powerful closed-source LLMs for detection.
@@ 3.1 Auto-Detection of Overthinking
As illustrated in Figure 1, we observe a similar phenomenon in R1-7B: the model tends to engage in excessive reflection after completing reasoning and obtaining the correct answer, leading to overthinking. In preliminary experiments, we find that weaker LLMs struggle to extract the boundaries of each reflection. However, we note that the conclusion of both reasoning and reflection is often a restatement of the answer. Therefore, we prompt LLMs to determine if each part of the response contains the correct answer, which is easier for LLMs. The thought process preceding the first occurrence of the correct answer is considered effective reasoning, while each subsequent appearance is an additional reflection, i.e., overthinking.
@@
Given inconsistent formatting and incidental early mentions of the answer in reasoning, regex-based extraction is unreliable. Thus, we employ Qwen2.5-32B-Instruct (Qwen-32B; Yang et al., 2024a) to identify these positions within the “think” part of LRMs. Specifically, we segment the think part into smaller chunks and provide the question along with these chunks, prompting the model to determine whether each chunk contains the correct answer. To ensure accuracy, we apply a filtering step: chunks containing the answer are verified again one by one with Qwen-32B, and only those confirmed as positive in both detections are considered correct. We design prompts with and without the inclusion of the gold answer as input, detailed in Appendix A.
@@
The tokens after the first correct answer are identified as overthinking and subsequently removed. Following this, we generate the revision by forcibly terminating the “think” part and compelling the LRM to continue generating the final answer, prefixed with ”**Final Answer:**”, which yields a revision without overthinking. For verification, the revision is deemed correct if the LRM can accurately generate the final answer. Finally, to prevent excessively long generations from skewing the average length, we limit the generation budget to 16k tokens.
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
@@
Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
Fixed Trunc (90) 89.39 80.51 91.40 83.51 81.30 83.20 85.00 85.68 44.17 88.72 78.25 84.32
Fixed Trunc (80) 88.25 72.47 89.60 75.52 79.22 75.22 78.44 77.67 37.08 80.09 74.52 76.19
Fixed Trunc (70) 87.26 64.73 86.00 66.76 75.58 66.61 73.12 68.99 33.33 71.02 71.06 67.62
Model Revise 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 50.83 93.59 80.61 75.65
  + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 51.67 92.31 79.82 66.32
Table 1: Performance and efficiency after model revision. “Original” denotes the R1-7B performance before revision. “Acc” represents the response accuracy, and “TR” represents the efficiency, calculated as its token cost divided by that of “Original”. Therefore, the TR in the first row is 100, and its subscript indicates its token cost. “Average” is the macro-average across datasets. The best and second-best results in the “Average” column are marked bold and underline, respectively. Considering the trade-off between TR and Acc, for TR, only methods whose accuracy does not decrease are bolded or underlined. Abbreviations of the methods are defined in §3.2.
@@ 3.2 Efficiency of Auto-Detection
We first use R1-7B to answer each question, and then apply the above response revision method from the previous section to truncate the generated reasoning path and complete it with R1-7B. We evaluate the correctness of response revision by the accuracy of the final answers after completion. “Model Revise” and “+Gold” represent revision using our method without and with the gold answer as input, respectively. Our baseline for comparison is “Fixed Trunc (nn)”, which denotes fixed truncation of the original generated response, retaining $n\%$ of the thinking tokens and truncating the response at the nearest newline. Subsequently, similar to response revision, the “think” part is terminated, and the LRM is compelled to generate the final answer. Results before and after revision are in Table 1. The evaluated math datasets are introduced in Appendix B.2, and are ordered from left to right with increasing difficulty.
@@
Without the gold answer as input, we can automatically remove 24% of tokens without performance degradation. When provided, 34% of tokens are removed with a minor performance decrease. However, fixed truncation shows a considerable performance decline with increasing truncation ratios. These results demonstrate the effectiveness of our detection method and underscore the severity of the overthinking problem.
Greater overthinking on easier problems.
Across three easier and two harder datasets, overthinking tokens our method detects decrease progressively, at 37% and 17%, respectively. This indicates that overthinking is more prevalent in easier datasets and less prevalent in more challenging ones.
@@ 4 Reflection-Aware Online Reinforcement Learning
To maintain model performance while reducing inference costs, a fundamental method is to employ online RL with a length reward4.1), which ensures distributional alignment between the data and the model. Building upon this, we make improvements from two perspectives. First, based on the detection method in §3, we introduce a reflection model that provides revisions online (§4.2). This not only augments the data but also provides shorter, non-overthinking paths that serve as positive examples for training, which are lacking in parallel sampling. Second, we refine the reward by including a reflection reward to penalize non-reflective behavior (§4.3), which would otherwise lead to a significant performance drop. The workflow is illustrated in Figure 2.